# Syntax-Guided Procedural Synthesis of Molecules

### v1

@@ 4.2.2 Synthesizable Molecule Design
Table 3: (Left) AutoDock Vina scores against DRD3 and MPro, limited to 5000 Oracle calls. For ZINC (Screening) we use numbers from TDC’s DRD3 Leaderboard. For SynNet, we report both their paper’s numbers and our reproduced runs. (Right) We report the top 3 binders for MPro for the real-world case study in App. H.
Top 1 Top 10 Top 100
Method Target Oracle Calls Score SA AUC Score SA AUC Score SA AUC
ZINC DRD3 N/A 12.8 -- -- 12.59 -- -- 12.08 -- --
Synnet 5000 10.8 2.589 10.092 10.3 2.592 9.547 9.197 2.355 8.374
Synnet (paper) 5000 12.3 2.801 -- 12.02 -- -- 11.133 -- --
Ours (top k) 5000 13.7 1.908 12.702 13.01 2.128 11.905 12.126 2.175 10.862
Synnet MPro 5000 8.3 2.821 8.015 8.09 2.267 7.602 7.46 2.249 6.816
Ours (top k) 5000 9.9 2.267 9.478 9.54 2.595 9.009 9.024 2.498 8.288
Method (#Oracle calls) 1st 2nd 3rd
SynNet (5000) 8.3 8.3 8.2
SynNet (Source: Paper) 10.5 9.3 9.3
Ours (5000) 9.9 9.7 9.7
Ours (10000) 10.8 10.7 10.6
Dataset Method RR $\uparrow$ Avg. Sim. $\uparrow$ SA Score $\downarrow$ Diversity $\uparrow$
$k=1$ $k=3$ $k=5$ $k=1$ $k=3$ $k=5$ $k=3$ $k=5$
Train Set Ours-reverse (Skeleton) 79.30% 0.923 0.632 0.569 3.072 2.795 2.716 0.615 0.657
Ours (Skeleton) 88.10% 0.958 0.704 0.626 3.099 2.928 2.852 0.532 0.615
Test Set SynNet 46.30% 0.766 0.622 0.566 3.108 3.057 3.035 0.525 0.584
Ours-reverse (Skeleton) 40.80% 0.749 0.548 0.487 2.97 2.743 2.659 0.64 0.685
Ours (Skeleton) 52.30% 0.799 0.588 0.548 3.075 2.895 2.856 0.609 0.653
1st 2nd 3rd $k=10$ $k=100$ Diversity
@@
Oracle Method $k=1$ $k=10$ $k=100$ Seeds All
QED SynNet 0.948 0.948 0.947 $0.673\pm 0.289$ $0.946\pm 0.001$
SynNet + BO 0.948 0.944 0.933 $0.622\pm 0.270$ $0.931\pm 0.007$
Ours (edits) 0.948 0.948 0.947 $0.391\pm 0.252$ $\textbf{0.947}\pm\textbf{0.000}$
@@
Ours (edits) 0.98 0.967 0.944 $0.074\pm 0.055$ $\textbf{0.941}\pm\textbf{0.012}$
@@
SynNet + BO 0.310 0.241 0.143 $0.006\pm 0.012$ $0.134\pm 0.039$
Ours (edits) 0.88 0.862 0.800 $0.059\pm 0.053$ $\textbf{0.792}\pm\textbf{0.030}$
@@
Ours (edits) 1.000 1.000 1.000 $0.024\pm 0.056$ $\textbf{1.000}\pm\textbf{0.000}$
Table 4: (Top left) Ablation 1: Top-down vs Bottom-up (Ours-reverse) Topological Order Decoding; (Bottom left) Ablation 2: Sibling pool generation strategies (edits: mutate skeleton, top $k$: top $k$ skeletons predicted from $Q_{\Theta}$, flips: don’t consider skeleton, flip random bit in the fingerprint) on JNK3 (Right) Ablation 3: SynNet with BO acquisition over sibling pool, generated via top $k$ beams
In Table 2, we see our method outperforming all synthesis-based methods on average across the 13 TDC Oracles for all considered metrics – average score, SA score, and AUC. Surprisingly, our method stays competitive with the SOTA string and graph methods in terms of Average Score (ranking 4th) but being considerably more sample-efficient at finding the top molecules (ranking 1st for Top 1/10 AUC). We see evidence that a synthesizability-constrained design space does not sacrifice end performance when reaping benefits of enhanced synthetic accessibility and sample efficiency.
The AutoDock Vina scores reflect our method’s strength in real-world ligand design tasks. Our best binders against $M^{\text{pro}}$ in Table 3 are significantly better than nearly all known inhibitors from virtual screening or literature [23, 70] ([70] reports a best score of -8.5). Our best binders against DRD3 also rank us 3rd on the TDC Leaderboard (as of Aug. 2024). We present additional analysis of the best binders for our two Docking targets in App H.
@@ 4.3 Ablation Studies
In this section, we analyze findings from 3 carefully designed ablation studies (numbered 1-3 in Table 4) to justify the key design decisions that differentiate our method from the predecessor SynNet as well as other synthesis-based methods that have similar modules. We leave a closer investigation of the structure-property relationship to App B.
4.3.1 Top-down vs Bottom-up Topological Order Decoding
@@
Our syntax tree is decoded top-down once the syntactic skeleton is fixed, but it can be argued a bottom-up decoding aligns better with reality and enables pruning, as done by all baselines that serialize the construction of synthesis trees ([2, 3, 21]) and assume the intermediate product satisfies the Markov property. We argue this formulation is ill-defined for early steps, e.g. the model has to predict the first building block to use given only knowledge of the target molecule. This is extremely difficult with orders of magnitude more building blocks than reactions (in our case, 147505 vs 91). Our method resolves this by reformulating the (Markov) state as partial syntax trees, where holes are reactions and building blocks left to predict. This state captures the horizon structure, so we can learn tailored policies for the fixed horizon. We reintroduce the inductive bias of retrosynthetic analysis to procedural synthesis. Conditioned on the syntax (skeleton), we argue a top-down filling order outperforms bottom-up; it is easier to reason backwards from the specification (target molecule) which reactions lead to the product. One may argue a pure bottom-up construction compensates by pruning reactions that don’t match any current precursors. To weigh how much this compensates, we perform an additional Ablation 1: instead of the MDP enforcing we fill in a skeleton top-down, we fill the skeleton from the bottom-up (i.e. a node can only be filled if its children are filled already). We retrain the model by pretraining on inverted masks, and decode by following every possible “reversed” topological order (i.e. topological order of the skeleton with edges reversed). Ablation 1 results show this cannot reconstruct the training data as well and cannot generalize. Conditioned on a syntactic template, we can conclude top-down decoding constrains the search significantly more, even if not able to prune on-the-fly.
4.3.2 Sibling Generation Strategy in Molecule Design
@@
Our analog generation capability is demonstrated in 4.2.1, but it’s not clear how or why it translates to better performance when used as an offspring generator within molecule optimization. Our surrogate takes as input a fingerprint and outputs multiple synthesizable analogs. A good surrogate should output offspring(s) that balance local neighborhood exploration (this Ablation) with global exploitation (Ablation 3). SynNet does the former by mutating the fingerprint directly, whereas the key insight of our syntax-guided method is to mutate the syntactic skeleton instead. Our method achieves this via editing mutations, but it’s not clear whether this is superior to the top recognition strategy employed for the analog generation task ((Skeleton) in Table 1). Table 4 suggests edit-based mutations are superior to the top recognition strategy used for analog generation and the trivial strategy of ignoring the skeleton and flipping individual bits to obtain siblings. This suggests the sibling pool related by edits to the skeleton is a better way to preserve a locality bias within the GA. The simultaneous increase to population diversity and average scores for $k=10,100$ also suggests the same symbiotic relationship between diversity and similarity in analog generation is also the key enabler to better GA optimization.
4.3.3 Comparison against SynNet with Sibling Generation
Our GA benefits from an inner-loop sibling acquisition within the crossover operation, acquiring the best sibling to expend an Oracle call on. It can be argued this extra mechanism is why our method gets better results and makes for an unfair comparison with SynNet, or that this is a method-agnostic hack to improve GA performance. However, we argue the performance gains of this mechanism is unlocked by our syntactic approach to generating a sibling pool. To prove this, we endow SynNet with a similar mechanism in its crossover operation, where we allow it to generate an offspring pool using the top beams then apply a BO acquisition step on top. However, the results not only didn’t improve, but actually downgraded the performance. We hypothesize the reason is SynNet’s optimization trajectory is not helped, but derailed by the additional variation to its sibling pool, reducing local movements within the output space that a syntactic editing approach naturally preserves.
@@ 5 Discussion & Conclusion
We reconceptualize the design of synthesis pathways through the lens of program synthesis. Drawing on ideas from syntax-guided synthesis, we introduce a bilevel framework that decouples the syntactical skeleton of a synthetic tree from its chemical semantics. We propose learning algorithms that amortizes over the search for the correct program semantics by fully utilizing the tree horizon structure imposed by the syntactic template. Our results demonstrate this approach’s advantages across the board on key evaluation metrics of conditional analog generation and showcase the method’s capability on challenging synthesizable de novo design problems. Presented with an more expressive design space $\mathcal{T}\times\mathcal{X}$, our method makes the most of the unique opportunities by decoupling the syntactic dimension to navigate a rich, joint design space. Our empirical results demonstrate our algorithms are a step towards capturing the full modeling potential of this design space.
Our bilevel framework not only integrates design and synthesis into a single workflow, significantly reducing the cycle time of traditional molecule discovery, but also invites experts into the loop. Our framework offers a degree of control over the resources involved in the synthesis process. Notably, syntactic templates give users a degree of control over the amount of resources required to execute synthesis and biases the design space towards simpler solutions, making it particularly promising for autonomous synthesis platforms. Combining our workflow with practical deployment, where resource optimization is critical, is an exciting avenue of work [10]. Our framework lays the foundation for an exciting vision: an interactive system that elicits domain expertise in the form of syntactic templates to expedite the procedural synthesis of new molecules. Similar to how syntax-guided program synthesis are meant to involve a programmer in-the-loop, we want to enable the expert to play a role in guiding the optimization towards better solutions.
@@ Appendix A Syntactic Templates
Syntactic templates form the essential ingredients for syntax-guided synthesis, as they significantly reduce the number of possible programs. In practice, syntactical templates are provided by users who operate with real-world constraints or experts who can help narrow the search space to desirable templates. The exact criteria for selecting templates are problem-dependent. To prove our concept in a more generalizable workflow, we bootstrap our set of syntactic templates $\hat{\mathcal{T^{\prime}}}$ in a data-driven way by obtaining the syntactic templates present in $D_{\text{train}}$. We then simulate real-world constraints by setting $\hat{\mathcal{T^{\prime}}_{k}}\leftarrow\{T^{\prime}\in\hat{\mathcal{T^{\prime}}}\mid\text{$T^{\prime}$ has at most $k$ reactions}\}$ and optimize within the design space $\hat{\mathcal{T^{\prime}}_{k}}$. We tabulate summary statistics in 5 for the number of unique syntactic templates and the number of topological orders. We see that the empirical distribution is biased towards simpler syntactic templates, which reflects real-world constraints and is a key enabler of our amortized approach. We train $(\Theta,\Phi,\Omega)_{k}$ for $k=3,4,5,6$ on $D_{\text{pretrain}}$. For samples in $D_{\text{pretrain}}$ with more than $k>6$ reactions, we snap it to the closest $T^{\prime}\in\hat{\mathcal{T^{\prime}}}$ according to the Tree Edit Distance. We find $k=3$ using full topological decoding (illustration in 2) is best for Synthesizable Analog Generation and $k=4$ with random sampling of the decoding beams is a good compromise between accuracy and efficiency for Synthesizable Molecule Design. We also note that the number of unique templates grows sub-exponentially, and in fact the number of templates for a fixed number of reactions starts diminishing for $k>6$. To make sure this does not cause issues, we ensured there is still sufficient coverage to formulate a Markov Chain on $\hat{\mathcal{T^{\prime}}_{k}}$, which is crucial for our bilevel algorithms. For example, 4 visualizes the empirical proposal distribution $J(T_{1},T_{2}),\forall T_{1},T_{2}\in\hat{\mathcal{T^{\prime}}_{4}}\times\hat{\mathcal{T^{\prime}}_{4}}$. Importantly, key hyperparameters like $\beta$ and $n_{\text{edits}}$ enable control over exploration vs exploitation.
Figure 4: We adopt the Tree Edit Distance as the dist function. We see that $\hat{\mathcal{T^{\prime}}_{4}}$ has sufficient transition coverage for bootstrapping our space of syntactic templates.
No. of Reactions No. Templates No. Topological Orders (Max, Mean, Std)
1 2 2, 1.5, 0.5
2 6 8, 4.17, 2.79
3 22 80, 19.59, 20.55
4 83 896, 152.02, 215.53
5 209 19200, 2506.25, 3705.77
No. of Reactions No. Templates
@@
Figure 5: (a) Summary statistics of the number of syntactic templates and possible topological decoding node orders for $k=1,2,\ldots,5$; (b) Summary statistics for only the number of syntactic templates since enumerating all topological sorts becomes intractable
Appendix B Syntax Tree Recognition

### v2

@@ 4.2 Results on Synthesizable Analog Generation
Table 3: AutoDock Vina scores against DRD3 and Mpro, limited to 5000 oracle calls. For ZINC (Screening), we use numbers from TDC’s DRD3 Leaderboard, and for SynNet, we report both their paper’s numbers (*) and our reproduced results. We also report the top 3 binders for Mpro for the real-world case study in App. I.
Score $\uparrow$ AUC $\uparrow$ SA $\downarrow$
Target Method #Calls 1st 2nd 3rd Top-10 Top-100 Top-10 Top-10
DRD3 ZINC -- 12.8 -- -- 12.59 12.08 -- --
SynNet* 5000 12.3 -- -- 12.02 11.13 -- --
SynNet 5000 10.8 10.4 10.3 10.30 9.20 9.55 2.59
Ours 5000 13.7 13.1 13.1 13.01 12.13 11.91 2.13
Mpro SynNet 5000 8.3 8.3 8.2 8.09 7.46 7.60 2.27
Ours 5000 9.9 9.7 9.7 9.54 9.02 9.01 2.59
SynNet* -- 10.5 9.3 9.3 -- -- -- --
Ours 10000 10.8 10.7 10.6 -- -- -- --
Table 4: (Left) Ablation of sibling pool generation strategies on JNK3: (edits) mutates the syntax, ($\tau$) uses the top skeletons predicted from $\tau{}$, and (flips) doesn’t consider skeleton and instead flips random bits in the fingerprint. (Right) Ablation of SynNet with our Bayesian optimization (BO) acquisition over a sibling pool generated by beam search. Seeds and All are the average scores of the initial and final populations.
1st 2nd 3rd Top-10 Top-100 Diversity
@@
Oracle Method Top-1 Top-10 Top-100 Seeds All
@@
Ours 0.98 0.967 0.944 $0.074\pm 0.055$ $\textbf{0.941}\pm\textbf{0.012}$
@@
SynNet + BO 0.31 0.241 0.143 $0.006\pm 0.012$ $0.134\pm 0.039$
Ours 0.88 0.862 0.800 $0.059\pm 0.053$ $\textbf{0.792}\pm\textbf{0.030}$
@@
Ours 1.000 1.000 1.000 $0.024\pm 0.056$ $\textbf{1.000}\pm\textbf{0.000}$
4.3 Results on Synthesizable Molecule Design
In Table 2, we see our method outperforming all synthesis-based methods on average across the 13 TDC oracles for all considered metrics – average score, AUC, and SA score. Surprisingly, our method stays competitive with the state-of-the-art string and graph methods in terms of average score (ranking 4th) but being considerably more sample-efficient at finding the top molecules (ranking 1st for top-1/10 AUC). We see evidence that a synthesizability-constrained design space does not sacrifice end performance when reaping benefits of enhanced synthetic accessibility and sample efficiency.
The AutoDock Vina scores reflect our method’s strength in real-world ligand design tasks. Our best binders against Mpro in Table 3 are significantly better than nearly all known inhibitors from virtual screening or literature (Ghahremanpour et al., 2020; Zhang et al., 2021). For example, Zhang et al. (2021) report a best score of $-8.5$. Our best binders against DRD3 also rank us 3rd on the TDC Leaderboard (as of Aug. 2024). We present additional analysis of the best binders for our two docking targets in App. I.
@@ 4.4 Ablation Studies
In Tables 1 and 4, we analyze findings from four carefully designed ablation studies (1 in 4.4, 2 & 3 in 4.4, 4 in 4.4) to justify the key design decisions that differentiate our method from the predecessor SynNet as well as other synthesis-based methods that have similar modules.
4.4.1 Top-down vs Bottom-up Decoding.
@@
Should we decode top-down or bottom-up? This separation exists between retrosynthesis vs molecular design. Baselines that serialize the construction of synthesis trees (Bradshaw et al., 2019; 2020; Gao et al., 2021) adopt the latter. Given only knowledge of the target molecule, the model has to predict the first building block, the first reaction, and so on. Our empirical findings confirm observations by Gao et al. (2021) that the first few steps are difficult, due to inherent ambiguity – multiple valid choices exist. Our method resolves this bottleneck by reformulating the (Markov) state as partial syntax trees, where holes are reactions and building blocks left to predict. This state captures the horizon structure, so we can learn tailored policies for the fixed horizon. We reintroduce the inductive bias of retrosynthetic analysis to procedural synthesis. Conditioned on the syntax (skeleton), we show a top-down filling order outperforms bottom-up, with two intuitive explanations: (1) there are orders of magnitude less reactions than building blocks (91 vs. 147,505) and (2) it is easier to reason backwards from the specification (target molecule) which reactions lead to the product. To demonstrate these factors compensate for any gains from bottom-up pruning of applicable reactions, we perform an additional ablation in Table 1. Instead of the MDP enforcing we fill in a skeleton top-down, we fill the skeleton from the bottom-up. We retrain the model by pretraining on inverted masks, and decode by following every possible topological order of the skeleton with edges reversed. The results show this cannot reconstruct the training data as well and struggles on generalization.
4.4.2 How Analog Generation Capabilities Translate to Design Capabilities.
@@
Why does our superior analog generation capability translate to better performance when used as an offspring generator within molecule optimization? An ideal surrogate takes as input a fingerprint and outputs multiple synthesizable analogs, creating diverse offspring(s) that balance local neighborhood exploration with global exploitation. SynNet does the former by mutating the fingerprint directly, whereas the key insight of our syntax-guided method is to mutate the syntactic skeleton instead, doing so via editing mutations. Table 4 (Left) shows edit-based mutations are superior to the top recognition strategy used for analog generation ((Skeleton) in Table 1) and the trivial strategy of ignoring the skeleton and flipping individual bits to obtain siblings. This suggests edits to the skeleton better preserves the locality bias within the GA. Higher population diversity and average scores for $k\in\{10,100\}$ suggest the same symbiotic relationship between diversity and similarity in analog generation is also the key enabler to better GA optimization. Our GA benefits from an inner loop sibling acquisition within the crossover operation, acquiring the highest expected improvement sibling to expend an oracle call on. It can be argued this extra mechanism is why our method gets better results and makes for an unfair comparison with SynNet, or that this is a method-agnostic hack to improve GA performance. In Table 4 (Right), we show SynNet endowed with a similar mechanism in its crossover operation (generate an offspring pool using the top beams then apply a BO acquisition step on top) didn’t improve, but actually downgraded the performance. We hypothesize that SynNet’s optimization trajectory is derailed by the additional variation to its sibling pool, reducing local movements within the output space that a syntactic editing approach naturally preserves. Thus, we believe the performance gains of this mechanism is unlocked by our syntax-driven approach.
4.4.3 Extrapolation to Unseen Templates.
How dependent is our framework on the initial set of syntactic templates? Can SynthesisNet generalize to more templates? Our ablation study in App. B evaluate whether the model can extrapolate to new template classes. We investigate: 1) how well SynthesisNet extrapolates to programs whose structural template was unseen during training, 2) how well the framework can incorporate new templates at test time, and 3) how robust overall task performance is to them. To answer these questions, we hold out $\approx 25\%$ of $\hat{\mathcal{T}}$ to be the test set, remove all programs belonging to those templates, retrain, and analyze changes in task performances, both holistic results and results specific to the held-out templates. Our results reveal minor performance drop, and in some instances, improved results along some metrics. Our analysis reveals the surrogate model is robust and unsensitive to missing templates, and the framework can also incorporate unseen templates in the online phases of the downstream tasks. We encourage future works to study the exact scaling laws between templates and performance, and we hope our initial findings unlock new directions for scaling up our solutions to achieve greater coverage and impact.
@@ 5 Discussion & Conclusion
We reconceptualize synthesis pathway design using a program synthesis approach, introducing a bi-level framework that separates the syntactical skeleton of a synthetic tree from its chemical semantics. Our learning algorithms leverage the tree horizon structure, improving performance on key metrics of analog generation and de novo design. By decoupling syntax and semantics, we effectively navigate a rich design space, integrating design and synthesis into a single workflow that reduces discovery cycle time. Our framework offers control over synthesis resources and biases towards simpler solutions, with the exciting prospect of integration with autonomous synthesis platforms (Coley et al., 2019).
@@ Appendix A Syntactic Templates
Syntactic templates form the essential ingredients for syntax-guided synthesis, as they significantly reduce the number of possible programs. In practice, syntactical templates are provided by users who operate with real-world constraints or experts who can help narrow the search space to desirable templates. The exact criteria for selecting templates are problem-dependent. To prove our concept in a more generalizable workflow, we bootstrap our set of syntactic templates $\hat{\mathcal{T}}$ in a data-driven way by obtaining the syntactic templates present in the training set. We then simulate real-world constraints by setting $\hat{\mathcal{T}_{k}}\leftarrow\{T\in\hat{\mathcal{T}}\mid\text{$T$ has at most $k$ internal nodes}\}$ and optimize within the induced design space $\hat{\partial{\mathcal{P}}}_{k}$ (Section 3.3.2). We tabulate summary statistics in Table 5 for the number of unique syntactic templates and the number of topological orders. We see that the empirical distribution is biased towards simpler syntactic templates, which reflects real-world constraints and is a key enabler of our amortized approach. We train the parameters $(\Theta,\Phi,\Omega)_{k}$ of our policies $(\tau,\pi_{{\mathcal{R}}},\pi_{{\mathcal{B}}})$, respectively, for $k=3,4,5,6$ on our (pre)training dataset ${\mathcal{D}}$. For samples in ${\mathcal{D}}$ with more than $k>6$ reactions, we snap it to the closest $T\in\hat{\mathcal{T}_{k}}$ according to the tree edit distance. We find $k=3$ using full topological decoding (illustration in Figure 3) is best for Synthesizable Analog Generation and $k=4$ with random sampling of the decoding beams is a good compromise between accuracy and efficiency for Synthesizable Molecule Design. We also note that the number of unique templates grows sub-exponentially, and in fact the number of templates for a fixed number of reactions starts diminishing for $k>6$. To make sure this does not cause issues, we ensured there is still sufficient coverage to formulate a Markov Chain on $\hat{\mathcal{T}}_{k}$, which is crucial for our bilevel algorithms. For example, Figure 4 visualizes the empirical proposal distribution $J(T_{1},T_{2}),\forall T_{1},T_{2}\in\hat{\mathcal{T}}_{4}\times\hat{\mathcal{T}}_{4}$. Importantly, key hyperparameters like $\beta$ and $n_{\text{edits}}$ enable control over exploration vs exploitation.
Figure 4: We adopt the tree edit distance as the dist function. We see that $\hat{\mathcal{T}}_{4}$ has sufficient transition coverage for bootstrapping our space of syntactic templates.
# Rxns $|\hat{\mathcal{T}}_{k\setminus k-1}|$ ($|\mathcal{T}_{k\setminus{k-1}}|$) # Topo. Orders (Max, Mean, Std)${}_{\hat{\mathcal{T}}_{k\setminus{k-1}}}$
1 2 (2) 2, 1.5, 0.5
2 6 (6) 8, 4.17, 2.79
3 22 (22) 80, 19.59, 20.55
4 83 (90) 896, 152.02, 215.53
5 209 (394) 19200, 2506.25, 3705.77
# Rxns $|\hat{\mathcal{T}}_{k\setminus k-1}|$
@@
# Rxns # Topo. Masks ()${}_{\hat{\mathcal{T}}_{k\setminus k-1}}$ Topo. Masks ()${}_{{\mathcal{T}}_{k\setminus k-1}}$
1 5, 4, 1 5, 4, 1
2 11, 7.67, 2.56 11, 7.67, 2.56
3 26, 14.36, 5.86 26, 14.36, 5.86
4 56, 27.99, 12.47 56, 26.73, 12.78
5 131, 65.07, 26.36 131, 49.74, 27.09
6 287, 165.12, 61.43 287, 92.67, 56.29
@@
Figure 5: (a) Summary statistics of the number of syntactic templates (both empirical and theoretically possible) and possible topological decoding node orders for $k=1,2,\ldots,5$; (b) Summary statistics for only the number of syntactic templates since enumerating all topological sorts becomes intractable; (c) Summary statistics for the number of topological masks (subset of nodes closed under parent(.))
Appendix B Extrapolation to Unseen Templates
B.1 Data Preparation
In this section, we investigate whether SynthesisNet can extrapolate to unseen templates, and effectively incorporate them for synthesizable analog generation and molecular design. We setup an ablation study as follows:
  1. 1.
    Reverse sort the templates by frequency using our dataset $\mathcal{D}_{0}$.
  2. 2.
    Collect every fourth template into a hold-out set $\mathcal{T}_{\text{test}}$ for $k=4$.
  3. 3.
    Construct $\mathcal{D}_{0}^{\prime}:=\{(P,B)\in\mathcal{D}_{0}\mid T_{P,B}\in\mathcal{T}_{4}\setminus\mathcal{T}_{\text{test}}\}$ where $T_{P,B}\in\mathcal{T}$ is the syntactical template of $(P,B)$.
  4. 4.
    Run Algo. 1 on $\mathcal{D}_{0}^{\prime}$ to obtain $\mathcal{D}^{\prime}$.
  5. 5.
    Train ablation policy networks $\{\pi_{\mathcal{B}}^{\prime},\pi_{\mathcal{R}}^{\prime}\}$ using $\mathcal{D}^{\prime}$.
  6. 6.
    Evaluate task performances with same $\tau$ as before.
We select hold-out templates in a frequency-stratified manner, ensuring the frequency distribution of $\hat{\mathcal{T}}_{\text{test}}$ is similar to that of $\hat{\mathcal{T}}$. Since smaller templates appear more frequently, the sizes of of templates are also indirectly stratified this way. Since we choose the least frequent from each consecutive group of 4 (Step 2), we note on average templates in $\hat{\mathcal{T}}_{\text{test}}$ tend to be slightly larger than $\hat{\mathcal{T}}$, so results for test templates may be lower in Table 5.
B.2 Results on Synthesizable Analog Generation
Table 5: Apart from swapping out the policy networks, we use the same experimental setup as Table 1. For fair comparison, we also retrained Ours with $k=4$ templates (whereas Table 1 used $k=3$).
Avg. Sim. $\uparrow$ SA $\downarrow$ Diversity $\uparrow$
Dataset Method RR $\uparrow$ Top-1 Top-3 Top-5 Top-1 Top-3 Top-5 Top-3 Top-5
Test Set Ours:EP ($\tau$) 52% 0.815 0.616 0.548 3.140 2.964 2.892 0.585 0.646
Ours ($\tau$) 56% 0.827 0.633 0.555 3.100 3.019 2.918 0.543 0.628
Ours:EP $\mathcal{T}_{\text{test}}$ ($\tau$) 20% 0.636 0.539 0.473 2.844 3.023 2.987 0.564 0.675
Ours $\mathcal{T}_{\text{test}}$ ($\tau$) 20% 0.626 0.552 0.493 3.360 3.135 3.070 0.542 0.634
ChEMBL Ours ($\tau$) 7.6% 0.531 0.443 0.396 2.544 2.510 2.460 0.675 0.727
Ours (MCMC) 9.2% 0.532 0.486 0.432 2.364 2.310 2.263 0.765 0.759
Ours:EP (MCMC) 8.5% 0.519 0.421 0.367 2.644 2.420 2.382 0.618 0.640
We evaluate the synthesizable analog task performances using ablation networks. We want to test whether the ablation networks integrate effectively with templates outside its structural support. Thus, we use the same $\tau$ as before for the $\tau$ experiments in Table 5, allowing access to the full template set but forcing $\{\pi_{\mathcal{B}}^{\prime},\pi_{\mathcal{R}}^{\prime}\}$ to extrapolate when performing inference over $\hat{\mathcal{T}}_{\text{test}}$.
The performance takes only a minor dip on the Test Set, compensating for slightly lower Avg. Sim. with higher Diversity and comparable SA. We further zoom in on the subset of $\mathcal{D}_{0}$ with structure among the held-out template classes. We emphasize this is very difficult for $\{\pi_{\mathcal{B}}^{\prime},\pi_{\mathcal{R}}^{\prime}\}$ to do, which has not seen any examples from those structural classes. It actually appears Ours:EP has the slight edge over Ours on the held-out template set, with slightly better SA and greater diversity. We attribute this to a regularization effect induced by removing (slightly, due to Step 2.) more complex program structures. There are still enough complex templates left that this does not harm performance, highlighting the robustness of our model in this setting. We believe the fact the ablation model can maintain comparable performance implies the following:
  • Structural Extrapolation: It is capable of inference of programs outside the structural support of its training distribution in this case.
  • Template Set Robustness: Our model is not very sensitive to the default size of the template set, since using only 75% of it already brings it close to diminishing returns.
The dip in performance is more noticeable for the predominantly unsynthesizable dataset ChEMBL, with lower metrics across the board. We suspect the reason is due to ChEMBL containing more complex molecules that require longer synthetic routes. This is also apparent from the lower analog diversity, suggesting training on more template variety helps.
We believe the difficulty of the task (ratio of synthesizable vs unsynthesizable molecules) can inform whether the method is sensitive to the template set it sees during training. Similar ablation studies to this one can highlight when additional resources should be allocated to expanding the training set and when it is sufficient to simply incorporate more templates at test time.
B.3 Results on Synthesizable Molecular Design
Figure 6: We select the first Oracle from each Table in App. H to compare Ours with Ours (EP). Aside from the ablation networks, we use the same experimental settings as Table 2.
[…]
Top 1 Top 10 Top 100 Top 1 Top 10 Top 100
category Oracle Calls Score SA AUC Score SA AUC Score SA AUC Oracle Calls Score SA AUC Score SA AUC Score SA AUC
Ours (EP) synthesis 6921 0.99 1.975 0.872 0.965 2.321 0.818 0.94 2.237 0.744 9050 0.4 4.12 0.358 0.344 4.434 0.305 0.301 4.394 0.257
Ours synthesis 4886 0.98 2.045 0.891 0.967 2.302 0.848 0.944 2.27 0.778 8303 0.4 3.353 0.371 0.342 4.161 0.305 0.298 4.256 0.252
Osimertinib MPO Perindopril MPO
Top 1 Top 10 Top 100 Top 1 Top 10 Top 100
category Oracle Calls Score SA AUC Score SA AUC Score SA AUC Oracle Calls Score SA AUC Score SA AUC Score SA AUC
Ours (EP) synthesis 10000 0.852 2.322 0.831 0.849 2.475 0.823 0.839 2.484 0.802 10000 0.626 3.382 0.562 0.598 3.375 0.54 0.56 3.349 0.509
Ours synthesis 10000 0.859 2.263 0.826 0.847 2.21 0.81 0.832 2.249 0.769 10000 0.622 3.338 0.547 0.591 3.378 0.524 0.558 3.137 0.485
We also evaluate the synthesizable molecular design performances using ablation networks. We want to evaluate whether the ablation models can guide the optimization trajectory as a surrogate generator of synthesizable analogs. We see comparable performances across all four Oracles in Table 14, and for each Ours (EP) having the edge on some metrics while Ours having the edge on other metrics. This suggests both are capable enough to serve as the inner subroutine of our bilevel genetic framework, although the models may have different biases on the kind of analogs it generates which affects the optimization trajectories. Since Ours (EP) may generate less structurally diverse analogs, it can converge in fewer Oracle calls, resulting in less Oracle calls and slightly higher AUCs. Meanwhile, Ours produce more diverse analogs, which enables the acquisition of higher confidence analogs. We see Ours to have the edge on SA across Median 1 (Top 1) and the MPO (Top 100) Oracles. This may be because higher confidence regions tend to be where the simpler molecules are, resulting in simpler analogs hence lower SA. Overall, the results show the robustness of our model.
Appendix C Syntax Tree Recognition